OpenStudy (thereisnousername):

Medal for best explanation: Let f(x) = −4x + 7 and g(x) = 10x − 6. Find f(g(x)).

OpenStudy (thereisnousername):

do I just add the two equations together?

satellite73 (satellite73):

no

satellite73 (satellite73):

\[f(g(x))=f(10x-6)\] is a start

satellite73 (satellite73):

then since \[f(\heartsuit)=-4\heartsuit+7\] you have \[f(10x-6)=-4(10x-6)+7\] then algebra

OpenStudy (thereisnousername):

sorry was writing, thanks for the help!

satellite73 (satellite73):

yw